OpenStudy (anonymous):

Harvey is 3 times as old as Jane. The sum of their ages is 48 years. Find the age of each. Jane is a0 years old. Harvey is a1 years old.

OpenStudy (texaschic101):

h + j = 48 h = 3j now sub 3j in for h 3j + j = 48 4j = 48 j = 12 h = 3j h = 3(12) h = 36 jane is 12 harvey is 36